# Do Current Language Models Support Code Intelligence for R Programming Language?

### v2

@@ 3.2.4. Parser
After this processing, the dataset was reduced to 1583 repositories, pertaining to 40,833 functions (i.e., code-comment pairs). The statistics of our collected dataset are in Table 1. The Base R dataset is shown with RBase, the RTidy represents the Tidy-verse data, and the RCombine is the original collected data, without separating Base and Tidy-verse. The other rows of the table demonstrate the dataset from CodeSearchNet. Follwoing the previous works (Husain et al., 2019), we split the RBase, RTidy, and RCombine with 8:1:1, as shown in Table 1. Note that we have ensured that there are no duplicate functions in the dataset.
@@ 3.3. Tasks
Code Summarization is a technique in software engineering that aims to automatically generate a concise and informative summary of the code  (Feng et al., 2020; Guo et al., 2020). Code summarization has the advantage of being fast, consistent, and reproducible, and can help developers easily understand the semantics of the source code (Zhu and Pan, 2019). This can be particularly useful in large and complex software projects where it can be difficult to keep track of the different components and understand how they interact with each other. In recent years, code summarization has gained significant attention in the software engineering community, with several research studies exploring the effectiveness of various summarization techniques and models (Devlin et al., 2019; Husain et al., 2019). However, previous research in code summarization has mainly focused on programming languages that are widely used in firms such as Java and Python  (Feng et al., 2020; Guo et al., 2020; Liu et al., 2019), leaving a gap in our understanding of code summarization for other programming languages, such as R. Therefore, we study the code summarization as one of the tasks. The dataset we collected includes both the methods and their natural language comments, written by developers. Therefore, we can use the collected dataset and it is ready to train the models for code summarization.
@@
Method Name Prediction or extreme code summarization refers to the task of predicting or generating appropriate names for methods (functions or procedures) in programming code (Allamanis et al., 2016). Given a code snippet, the model aims to generate a suitable name that accurately represents the intended behavior or functionality of the method. This task is important in software development as well as code comprehension because well-named methods can improve code readability, maintainability, and understandability (Allamanis et al., 2016). Method Name Prediction often involves training models on large-scale code repositories to learn patterns and common naming conventions used in programming. The models can then use this knowledge to generate plausible method names for given code contexts. This task has applications in various code-related tasks, such as code completion, code summarization, and code search, where generating informative method names can greatly assist developers in writing or understanding code more efficiently (Petukhov et al., 2022). It also has been studied in multiple previous research (Ahmed and Devanbu, 2022; Alon et al., 2019, 2018). Therefore, we chose this task in our study. The collected dataset is valid for the method name prediction task, as given the code without the method name, and the natural language comments, we can train and evaluate the models to generate the method name.
@@
Code Generation aims to generate a code snippet from a given natural language description. Ridnik et al. (2024) purposed AlphaCodium which performed well on code generation benchmark. The model is given as input a description written in natural language, and the goal is to generate the code that corresponds to the description. Developers often take code from blogs, posts, articles, and other websites and adapt it to their context. The code task generation task is important as it provides a better replacement solution by generating these codes for users based on natural language descriptions. We select code generation to assess the capabilities of the models for a different category of code intelligence task (natural language to Code task). As there is no code generation for R, we will follow the HumanEval (Chen et al., 2021) dataset, which is in Python. We use MultiPL-E (Cassano et al., 2023) to evaluate the LLM’s performance on R. MultiPL-E (Cassano et al., 2023) is a multi-language benchmark that translates the HumanEval dataset from Python into 18 other programming languages. We sample 20 completions per prompt at temperature 0.2 with top-p 0.9 following (Cassano et al., 2023). Note that typically 50 is needed to compute Pass@1 but Cassano et al. (2023) reports that 20 completions would give a stable estimate for the Pass@1 score. Note that due to a lack of training data for code generation in R, we will run the experiments in a zero-shot paradigm, meaning that the models will not be fine-tuned for this task. Instead, we only ask the model to generate code without showing an example (i.e., zero-shot).
@@ 3.4. Studied Code-PLMs
In this section, we will briefly discuss the chosen benchmarking pre-trained models. We choose these models based on their architecture, pre-training tasks, and performance, and we follow the recommendation of a recent empirical study for selecting them (Niu et al., 2023). Niu et al. (2023) conducted empirical studies on various code intelligence tasks and code-PLMs and provided recommendations for the model choices for each task. Accordingly, we choose CodeBERT, GraphCodeBERT, UniXcoder, and CodeT5 in our study. Note that at the time of this study, to the best of our knowledge, there is no research on evaluating the models for R language. Therefore, we follow the recommendations of (Niu et al., 2023) and choose the following models, in addition to conducting experiments with large language models, GPT-3.5-Turbo and CodeLlama.
CodeBERT (Feng et al., 2020) is a language model specifically designed for programming languages. CodeBERT’s architecture is based on the Transformer model, similar to the original GPT model, but with modifications to accommodate programming code. CodeBERT is pre-trained using two primary tasks: masked language modeling (MLM) and replaced token detection (RTD). For MLM, which has been shown to be effective in literature (Devlin et al., 2019), random tokens in the input code are masked, and the model is trained to predict the masked tokens. For RTD (Clark et al., 2020), the objective is to use both bimodal and unimodal data for training. CodeBERT has shown promising results in various programming language-related tasks, such as natural language code retrieval, and code summarization. It has achieved state-of-the-art performance at the time of publication in code-related benchmarks and competitions, demonstrating its effectiveness in understanding and generating code.
@@
UniXcoder (Guo et al., 2022) is a model designed for multilingual code-mixing scenarios, where programming code contains multiple languages. It aims to handle code written in multiple programming languages within the same codebase. UniXcoder employs a multi-modal architecture that combines the strengths of both Transformer models and recurrent neural networks (RNNs). It consists of a hybrid Transformer-RNN model, where the Transformer processes the textual code tokens, and the RNN handles the language-related features. UniXcoder utilizes pre-training tasks similar to CodeBERT to learn code-related representations. Additionally, it incorporates tasks specific to multilingual code-mixing scenarios, such as language identification and code-switching prediction. UniXcoder has demonstrated effectiveness in handling multilingual code-mixing scenarios and achieving high performance in code-related tasks involving multiple programming languages.
@@
CodeLlama-7B. CodeLlama-7B is a model from the Code Llama (Roziere et al., 2023) model series which is a large programming domain-oriented language model based on Llama 2 (Touvron et al., 2023), which allows for the direct generation or understanding of code using textual hints (Prompts). CodeLlama has the ability to perform code completion by generating up to 100k tokens (Roziere et al., 2023). Additionally, CodeLlama has the ability to generate zero-sample instructions for programming tasks. In addition, CodeLlama has zero-sample instruction adherence for programming tasks including code generation. This small-sample learning capability is worth noting when testing code to accomplish tasks in small resource languages like R. Roziere et al. (2023) claim that CodeLlama can be used for a wide range of programming tasks and is one of the best performing LLM at public programming tasks, making developers’ workflows faster and more efficient, and lowering the learning threshold for programming. Therefore, we select CodeLlama to explore the RQ5 answer.
@@
StarCoder2-base. In recent years, the field of code generation has advanced significantly with the development of Large Language Models (LLM) tailored for code-related tasks. One of the most prominent models is StarCoder (Li et al., 2023), along with its base variant, StarCoderBase, both part of the BigCode project [^10] https://huggingface.co/bigcode.StarCoder and StarCoder-base are models trained on permissively licensed data from GitHub, covering over 80 programming languages, Git commits, GitHub issues, and Jupyter notebooks. StarCoder-Base, which has approximately 15 billion parameters, was trained on 1 trillion tokens. StarCoder-Base was further fine-tuned on 35 billion Python tokens to create StarCoder, specifically optimized for Python-related tasks.
@@
StarCoderBase surpasses many existing open-source Code LLMs on popular programming benchmarks and competes with closed models like OpenAI’s code-cushman-001 (the Codex model behind early versions of GitHub Copilot). With a context length exceeding 8,000 tokens, StarCoder can handle large inputs, making it ideal for tasks like acting as a technical assistant, autocompleting code, modifying code based on instructions, or providing natural language explanations of code snippets. Given Starcoder’s performance and versatility, StarCoder models offer a strong foundation for community adoption and customization across various applications including code generation. Therefore, we use StarCoder-base in our experiments in RQ5.
@@
Codegen-2B-multi. We should note that, initially we intended to use CahtGPT as one of the LLMs for code generation. Since GPT-3.5 (chatgpt) is not open-source, we used CodeGen-2B-multi as an alternative for the code generation task. CodeGen-2B-multi is a large-scale transformer model trained on both natural language and multiple programming languages, allowing it to generate code from natural language descriptions. CodeGen-2B-multi is open-source, making it a valuable resource for research and development. The model is pre-trained on natural language datasets and fine-tuned on multilingual code datasets, enabling it to perform well across various programming languages for tasks like code completion and synthesis. Therefore, we used it for the experiments in RQ5.
@@ 3.5. Approach
RQ1: This RQ is motivated by the lack of research and understanding of the performance of Code-PLMs for the R programming language. To address this challenge, we compared the four Code-PLMs for each of the languages on the CodeXGLUE benchmark (which is the same data as CodeSearchNet for these tasks), and the R dataset. To achieve the comparison, we trained each of the models on each of the programming languages shown in the first six rows of Table 1 and also on the RCombine datasets separately, and then tested on the test split of each one. Therefore, we can evaluate them when trained and tested on each of the languages in a mono-lingual setting.
@@
RQ3: For this RQ, we use the RCombine dataset and we split the functions of the training and test set based on the repository they belong to. For the intra-project setting, we spited the training and test sets such that we have some functions from every repository in the training set and some of its functions in the test set. Note that we take care of not having duplicate functions in the training and test sets. For the cross-project setting, the repositories are split such that their functions only belong to the training or test set. Therefore, the model does not have functions from the same repository in the training and test set.
RQ4: As mentioned earlier in section 3, there are two different opinionated programming styles in R, the Base version, and Tidy-verse. The Base R is the fundamental R programming language, and it is the only dialect for a long time it can be more stable when compared with Tidy-verse (Wickham, 2014). Tidy-verse, on the other hand, is a collection of packages that work as an add-on to base R that provides extra functionality and makes the workflow and syntax more user-friendly. Despite the benefits and user groups of each, the syntax can be very different. Listing 1 provides an example of Tidy-verse and a Base R example is shown in Listing 2. Based on these differences, we hypothesize that separating the datasets might affect the obtained scores in a positive way.
@@
RQ5: To answer this research question, we run experiments using CodeT5 (Wang et al., 2021b) and UniXCoder (Guo et al., 2022), as well as ChatGPT (Brown et al., 2020), CodeLlama (Roziere et al., 2023) and StarCoder2 (Li et al., 2023). We exclude other models as they are not developed for code generation. For code generation, as stated earlier, the models will be prompted to generate code. We prompted ChatGPT (Brown et al., 2020) and CodeLlama in a zero-shot setting and compare its results with the full fine-tuned models. For code summarization and method name prediction tasks, we selected 100 samples from the test dataset (same as the ones used for human evaluation, described in Section 3.7).
@@ 3.6. Evaluation Metrics
For code summarization, following previous works (Feng et al., 2020; Wang et al., 2021b), we evaluate the results using smoothed BLEU-4 score (Lin and Och, 2004). The smoothed BLEU-4 score is a technique used in the field of machine translation to assess the quality of translations. It is an extension of the BLEU (Bilingual Evaluation Understudy) metric, which measures the similarity between a machine-generated translation and one or more reference translations. The BLEU score can be calculated using the following formula1:
@@
Statistical tests. Following the previous works (Ahmed and Devanbu, 2022), we apply the Wilcoxon signed-rank test to compare the results of the models for all the tasks.
@@ 3.7. Human Evaluation
Other than automatic evaluation metrics, we also conduct a human evaluation, following the protocol used in previous studies (Shahbazi and Fard, 2023; Sharma et al., 2022a). We randomly select $100$ samples from the test dataset for each task. Each sample will include code, ground truth values, and the generated output (code comments or method names) for all models. Two Computer Science graduate students, who are both familiar with R, conducted the evaluation. Each record was ranked by two reviewers and the average value were reported. Each sample was evaluated based on the following criteria, on a Likert scale of $1$ to $5$, where $1$ shows the lowest and $5$ shows the highest.
@@
  • Relevance (R), which evaluates the consistency of the provided details in the comments and compares it with the details in the code.
@@
For method name prediction, we considered the semantic similarity of the ground truth and the generated names. To avoid bias, we hided whether the text (comment or method name) is generated by a model or is written by a human. Moreover, the evaluators were educated before scoring. We provided examples and explanations for each criterion before they start evaluation.
@@ 3.8. Experimental Setup
All the experiments were run on the UBC Advanced Research Computing Sockeye and Digital Research Alliance of Canada. Each job requires the allocation of a compute node with 1 CPU core, 48GB of memory, and 2 NVIDIA Tesla V100 32GB GPUs with 32GB of memory for execution. For hyper-parameters used during the training phase, we followed the setting introduced in the original papers. We trained the models for 10 epochs with a learning rate of $0.00005$. The source_length is set to $256$ and the target_length is to $128$ for code summarization. For method name prediction, the target length is adjusted to $10$. The batch size is set to $32$ for CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020), $48$ for UnixCoder (Guo et al., 2022) and CodeT5 (Wang et al., 2021b) to perform a fair comparison to the original paper and the between R language and other programming languages. All experiments are run on Dell EMC R440 GPU compute nodes.

### v3

@@ 3.2.4. Parser
[…]
@@ 3.3. Tasks
Code Summarization is a technique in software engineering that aims to automatically generate a concise and informative summary of the code  (Feng et al., 2020; Guo et al., 2020). Code summarization has the advantage of being fast, consistent, and reproducible, and can help developers easily understand the semantics of the source code (Zhu and Pan, 2019). This can be particularly useful in large and complex software projects where it can be difficult to keep track of the different components and understand how they interact with each other. In recent years, code summarization has gained significant attention in the software engineering community, with several research studies exploring the effectiveness of various summarization techniques and models (Devlin et al., 2019; Husain et al., 2019). However, previous research in code summarization has mainly focused on programming languages that are widely used in firms, such as Java and Python  (Feng et al., 2020; Guo et al., 2020; Liu et al., 2019), leaving a gap in our understanding of code summarization for other programming languages, such as R. Therefore, we study code summarization as one of the tasks. The dataset we collected includes both the methods and their natural language comments, written by developers. Therefore, we can use the collected dataset, and it is ready to train the models for code summarization.
@@
Method Name Prediction or extreme code summarization refers to the task of predicting or generating appropriate names for methods (functions or procedures) in programming code (Allamanis et al., 2016). Given a code snippet, the model aims to generate a suitable name that accurately represents the intended behavior or functionality of the method. This task is important in software development as well as code comprehension because well-named methods can improve code readability, maintainability, and understandability (Allamanis et al., 2016). Method Name Prediction often involves training models on large-scale code repositories to learn patterns and common naming conventions used in programming. The models can then use this knowledge to generate plausible method names for given code contexts. This task has applications in various code-related tasks, such as code completion, code summarization, and code search, where generating informative method names can greatly assist developers in writing or understanding code more efficiently (Petukhov et al., 2022). It has also been studied in multiple previous research (Ahmed and Devanbu, 2022; Alon et al., 2019, 2018). Therefore, we chose this task in our study. The collected dataset is valid for the method name prediction task, as given the code without the method name, and the natural language comments, we can train and evaluate the models to generate the method name.
@@
Code Generation aims to generate a code snippet from a given natural language description. Ridnik et al. (2024) proposed AlphaCodium, which performed well on the code generation benchmark. The model is given as input a description written in natural language, and the goal is to generate the code that corresponds to the description. Developers often take code from blogs, posts, articles, and other websites and adapt it to their context. The code generation task is important as it provides a better replacement solution by generating code for users based on natural language descriptions. We select code generation to assess the capabilities of the models for a different category of code intelligence tasks (natural language to Code task). As there is no code generation for R, we will follow the HumanEval (Chen et al., 2021) dataset, which is in Python. We use MultiPL-E (Cassano et al., 2023) to evaluate the LLM’s performance on R. MultiPL-E (Cassano et al., 2023) is a multi-language benchmark that translates the HumanEval dataset from Python into 18 other programming languages. We sample 20 completions per prompt at temperature 0.2 with top-p 0.9 following (Cassano et al., 2023) for zero-shot samples. Note that typically 50 is needed to compute Pass@1, but Cassano et al. (2023) reports that 20 completions would give a stable estimate for the Pass@1 score.
We also run the expressions in the few-shot setting to compare different prompting techniques as well. We conducted one, three, five, and ten-shot experiments following the greedy approach (0 temperature) under three different prompting approaches: simple R, BM25 (Robertson and Zaragoza, 2009), and embedding-based using voyage-code-3 model. The details are discussed in Section 4.5.
@@ 3.4. Studied Code-PLMs
In this section, we will briefly discuss the chosen benchmarking pre-trained models. We choose these models based on their architecture, pre-training tasks, and performance, and we follow the recommendation of a recent empirical study for selecting them (Niu et al., 2023). Niu et al. (2023) conducted empirical studies on various code intelligence tasks and code-PLMs and provided recommendations for the model choices for each task. Accordingly, we choose CodeBERT, GraphCodeBERT, UniXCoder, and CodeT5 in our study. Note that at the time of this study, to the best of our knowledge, there is no research on evaluating the models for the R language. Therefore, we follow the recommendations of (Niu et al., 2023) and choose the following models, in addition to conducting experiments with large language models, GPT-3.5-Turbo and CodeLlama.
CodeBERT (Feng et al., 2020) is a language model specifically designed for programming languages. CodeBERT’s architecture is based on the Transformer model, similar to the original GPT model, but with modifications to accommodate programming code. CodeBERT is pre-trained using two primary tasks: masked language modeling (MLM) and replaced token detection (RTD). For MLM, which has been shown to be effective in literature (Devlin et al., 2019), random tokens in the input code are masked, and the model is trained to predict the masked tokens. For RTD (Clark et al., 2020), the objective is to use both bimodal and unimodal data for training. CodeBERT has shown promising results in various programming language-related tasks, such as natural language code retrieval and code summarization. It has achieved state-of-the-art performance at the time of publication in code-related benchmarks and competitions, demonstrating its effectiveness in understanding and generating code.
@@
UniXCoder (Guo et al., 2022) is a model designed for multilingual code-mixing scenarios, where programming code contains multiple languages. It aims to handle code written in multiple programming languages within the same codebase. UniXCoder employs a multimodal architecture that combines the strengths of both Transformer models and recurrent neural networks (RNNs). It consists of a hybrid Transformer-RNN model, where the Transformer processes the textual code tokens, and the RNN handles the language-related features. UniXCoder utilizes pre-training tasks similar to CodeBERT to learn code-related representations. Additionally, it incorporates tasks specific to multilingual code-mixing scenarios, such as language identification and code-switching prediction. UniXCoder has demonstrated effectiveness in handling multilingual code-mixing scenarios and achieving high performance in code-related tasks involving multiple programming languages.
@@
CodeLlama-7B. CodeLlama-7B is a model from the Code Llama (Roziere et al., 2023) model series, which is a large programming domain-oriented language model based on Llama 2 (Touvron et al., 2023), which allows for the direct generation or understanding of code using textual hints (Prompts). CodeLlama has the ability to perform code completion by generating up to 100k tokens (Roziere et al., 2023). Additionally, CodeLlama has the ability to generate zero-sample instructions for programming tasks. In addition, CodeLlama has zero-sample instruction adherence for programming tasks, including code generation. This small-sample learning capability is worth noting when testing code to accomplish tasks in small resource languages like R. Roziere et al. (2023) claim that CodeLlama can be used for a wide range of programming tasks and is one of the best performing LLM at public programming tasks, making developers’ workflows faster and more efficient, and lowering the learning threshold for programming. Therefore, we select CodeLlama to explore the RQ5 answer.
@@
StarCoder2-base. In recent years, the field of code generation has advanced significantly with the development of Large Language Models (LLMs) tailored for code-related tasks. One of the most prominent models is StarCoder (Li et al., 2023), along with its base variant, StarCoderBase, both part of the BigCode project [^11] https://huggingface.co/bigcode.StarCoder and StarCoder-base are models trained on permissively licensed data from GitHub, covering over 80 programming languages, Git commits, GitHub issues, and Jupyter notebooks. StarCoder-Base, which has approximately 15 billion parameters, was trained on 1 trillion tokens. StarCoder-Base was further fine-tuned on 35 billion Python tokens to create StarCoder, specifically optimized for Python-related tasks.
@@
StarCoderBase surpasses many existing open-source Code LLMs on popular programming benchmarks and competes with closed models like OpenAI’s code-cushman-001 (the Codex model behind early versions of GitHub Copilot). With a context length exceeding 8,000 tokens, StarCoder can handle large inputs, making it ideal for tasks like acting as a technical assistant, autocompleting code, modifying code based on instructions, or providing natural language explanations of code snippets. Given StarCoder’s performance and versatility, StarCoder models offer a strong foundation for community adoption and customization across various applications, including code generation. Therefore, we use StarCoder-base in our experiments in RQ5.
@@
Codegen-2B-multi. We should note that initially, we intended to use ChatGPT as one of the LLMs for code generation. Since GPT-3.5 (chatgpt) is not open-source, we used CodeGen-2B-multi (Nijkamp et al., 2022) as an alternative for the code generation task. CodeGen-2B-multi is a large-scale transformer model trained on both natural language and multiple programming languages, allowing it to generate code from natural language descriptions. CodeGen-2B-multi is open-source, making it a valuable resource for research and development. The model is pre-trained on natural language datasets and fine-tuned on multilingual code datasets, enabling it to perform well across various programming languages for tasks like code completion and synthesis. Therefore, we used it for the experiments in RQ5.
@@ 3.5. Approach
RQ1: This RQ is motivated by the lack of research and understanding of the performance of Code-PLMs for the R programming language. To address this challenge, we compared the four Code-PLMs for each of the languages on the CodeXGLUE benchmark (which is the same data as CodeSearchNet for these tasks) and the R dataset. To achieve the comparison, we trained each of the models on each of the programming languages shown in the first six rows of Table 1 and also on the RCombine datasets separately, and then tested on the test split of each one. Therefore, we can evaluate them when trained and tested on each of the languages in a mono-lingual setting.
@@
RQ3: For this RQ, we use the RCombine dataset, and we split the functions of the training and test sets based on the repository they belong to. For the intra-project setting, we split the training and test sets such that we have some functions from every repository in the training set and some of its functions in the test set. Note that we take care not to have duplicate functions in the training and test sets. For the cross-project setting, the repositories are split such that their functions only belong to the training or test set. Therefore, the model does not have functions from the same repository in the training and test sets.
RQ4: As mentioned earlier in section 3, there are two different opinionated programming styles in R, the Base version, and Tidy-verse. The Base R is the fundamental R programming language, and it has been the only dialect for a long time. Base R can be more stable when compared with Tidy-verse (Wickham, 2014). Tidy-verse, on the other hand, is a collection of packages that work as an add-on to base R that provides extra functionality and makes the workflow and syntax more user-friendly. Despite the benefits and user groups of each, the syntax can be very different. Listing 1 provides an example of Tidy-verse, and a Base R example is shown in Listing 2. Based on these differences, we hypothesize that separating the datasets might affect the obtained scores in a positive way.
@@
RQ5: To answer this research question, we run experiments using UniXCoder (Guo et al., 2022), as well as Codegen-2B (Nijkamp et al., 2022), CodeLlama (Roziere et al., 2023) and StarCoder2 (Li et al., 2023). We exclude other models as they are not developed for code generation. For code generation, as stated above, the models will be prompted to generate code. We prompted the LLMs in a zero-shot and few-shot setting and compare their results with the full fine-tuned models. For code summarization and method name prediction tasks, we selected 100 samples from the test dataset (same as the ones used for human evaluation, described in Section 3.7).
@@ 3.6. Evaluation Metrics
For code summarization, following previous works (Feng et al., 2020; Wang et al., 2021b), we evaluate the results using the smoothed BLEU-4 score (Lin and Och, 2004). The smoothed BLEU-4 score is a technique used in the field of machine translation to assess the quality of translations. It is an extension of the BLEU (Bilingual Evaluation Understudy) metric, which measures the similarity between a machine-generated translation and one or more reference translations. The BLEU score can be calculated using the following formula 1:
@@
Statistical tests: Following the previous works (Ahmed and Devanbu, 2022), we apply the Wilcoxon signed-rank test to compare the results of the models for all the tasks.
@@ 3.7. Human Evaluation
Other than automatic evaluation metrics, we also conduct a human evaluation, following the protocol used in previous studies (Shahbazi and Fard, 2023; Sharma et al., 2022a). We randomly select $100$ samples from the test dataset for each task. Each sample will include code, ground truth values, and the generated output (code comments or method names) for all models. Two Computer Science graduate students, who are both familiar with R, conducted the evaluation. Each record was ranked by two reviewers, and the average value was reported. Each sample was evaluated based on the following criteria, on a Likert scale of $1$ to $5$, where $1$ shows the lowest and $5$ shows the highest.
@@
  • Relevance (R), which evaluates the consistency of the provided details in the comments and compares them with the details in the code.
@@
For method name prediction, we considered the semantic similarity of the ground truth and the generated names. To avoid bias, we hide whether the text (comment or method name) is generated by a model or is written by a human. Moreover, the evaluators were educated before scoring. We provided examples and explanations for each criterion before they started the evaluation.
@@ 3.8. Experimental Setup
All experiments were run on the UBC Advanced Research Computing Sockeye and Digital Research Alliance of Canada. Each job required a compute node with 1 CPU core, 48  GB of system memory, and two NVIDIA Tesla V100 32 GB GPUs. We followed the hyper-parameter settings described in the original papers, training the models with a learning rate of 0.00005. The source_length was set to 256, and the target_length was 128 for code summarization. For method name prediction, the target_length was adjusted to 10. We employed a RandomSampler, ensuring that training samples—including those from different programming languages—were shuffled at each epoch to avoid biases in data ordering.
The batch size was fixed at 32 for CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020), and 48 for UniXCoder (Guo et al., 2022) and CodeT5 (Wang et al., 2021b), mirroring the original paper settings for a fair comparison across different programming languages. Finally, all experiments were executed on Dell EMC R440 GPU compute nodes to ensure consistent and reproducible results.